# The Cone Restriction: An old Approach Revisited

### v5

@@
Abstract
We consider Ou-Wang’s approach to cone restriction via polynomial partitioning. By restructuring their induction arguments into a recursive algorithm and applying the nested polynomial Wolff axioms, we refine the bounds on cone restriction estimate in higher dimensions.
1 Introduction
In this paper, we present a small improvement on cone restriction in higher dimensions. The Fourier restriction is one of the most significant open challenges in harmonic analysis, addresses the restriction of the Fourier transform to a compact manifold. Stein [16] conjectured that the restriction of Fourier transforms to a well-curved surface, such as sphere, paraboloids, or the cone studied in this paper, is a bounded $L^{p^{\prime}}\to L^{q^{\prime}}$ operator for some $p^{\prime}$ and $q^{\prime}$. The precise statement of cone restriction conjecture can be formulated as follows.
@@ 1 Introduction
there holds
$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{p^{\prime}},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
@@
$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}|\xi|)}f(\xi)\,d\xi\$
where $B^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that conjecture 1.1 is equivalent to
@@
Let $n\geq 3$. For all $p,q>0$ satisfy (1.2),
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
Conjecture 1.2 has so far been proven only in dimensions $n=3,4,5$ with contributions from Barcelo [1], Wolff [18], and Ou-Wang [15]. Prior to this work, the best-known range for higher dimensions was established by Ou-Wang [15], with results as follows.
@@
Let $n\geq 3$. For all $p,q>2$ satisfy
$q\geq p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$.
there holds
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
Theorem 1.4 ([15], Theorem 2).
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
@@
then
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
The Fourier restriction problem has been the focus of extensive study for decades. For hypersurfaces with non-vanishing Gaussian curvature, such as the sphere and paraboloid, restriction estimates are closely tied to several central problems in harmonic analysis, discrete geometry, and PDE, including the Bochner–Riesz conjecture, the Kakeya conjecture, and the Schrödinger maximal estimate. Furthermore, Fourier restriction has applications beyond harmonic analysis, impacting fields such as geometric measure theory (e.g., Falconer’s distance set conjecture) and number theory (e.g., counting solutions to Diophantine equations). However, sharp restriction estimates are known only for very few surfaces and dimensions. For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$ It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [14]), which highlights the potential applications of our results.
In this work, we show that:
@@
For $n\geq 3$,
$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,\epsilon}R^{\epsilon}\|f\|_{2}^{2/p}\|f\|_{\infty}^{1-2/p}$ (1.3)
whenever
$p>2+\lambda n^{-1}+O(n^{-2})$ (1.4)
where $\lambda=2.596$… and $B_{R}$ is an arbitrary ball in $\mathbb{R}^{n}$ of radius $R$.
By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [17]), one has
@@
Let $n\geq 3$. For all $q\geq p>2$ satisfy (1.4), there holds
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids. Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7]. In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, HRZ, 24], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [21, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
In this article, we refine the results presented in [15] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [11, 12]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [15] as a recursive process. Compared to the paraboloid case addressed in [11, 12], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor. We will provide a more detailed discussion on this topic in Section 3.
@@
[…]
Acknowledgements The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions. The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
2 Preliminaries
2.1 Notational Conventions
@@
In this paper, we work with Schwartz function $f$ or $g$ supported in $2\bar{B}^{n-1}\setminus B^{n-1}$. We write $x\in\mathbb{R}^{n}$ as $x=(x^{\prime},x_{n})$. The parameters $n,k,\epsilon$ are fixed. All implicit constants are allowed to depend on $n,k,\epsilon$. We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ as bellows:
$\delta_{0}:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We say $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$. Let #A\#A denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}$ denote an arbitrary ball in $\mathbb{R}^{n}$ of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$. We say $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$. We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$. We say $f$ is essentially supported in $\Omega$ if $f(x)\lesssim$ RapDec($R$) $\|f\|_{2}$. We say $Ef$ is essentially supported in $\Omega$ if $Ef(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
2.2 Wave Packet Decomposition
@@
The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
[…]
Figure 1: A sector of a truncated cone.
Given a function $f$ with supp $f\subseteq 2\bar{B}^{n-1}\setminus B^{n-1}$, one writes $f=\sum_{\theta\in\Theta[r]}f_{\theta}$ where $f_{\theta}=f\psi_{\theta}$ for $\psi_{\theta}$ a smooth partition of unity subordinate to the cover $\theta$. For each $\theta$, we break up $f_{\theta}$ as follows. First, we decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length  $r^{1/2+\delta}$, centered at $v\in r^{1/2+\delta}Z^{n-1}$. Now, we can write $\hat{f_{\theta}}=\sum_{v}\hat{f_{\theta}}\eta_{v}$ , where $\eta_{v}$ is a smooth partition of unity subordinate to this cover. Second, for each $v$, we decompose $\hat{f_{\theta}}\eta_{v}$ into even finer pieces according to $\theta$. Cover the cube centered at $v$ by parallel tine plates $P_{\theta,v}^{l}$ with radius  $r^{1/2+\delta}$, thickness  $r^{\delta/2}$, and normal direction $\xi_{\theta}$. Here $\xi_{\theta}$ is the point on the central line of a sector $\theta$ with $|\xi_{\theta}|=1$. Hence, one can write
$f=\sum_{\theta,v,l\in\mathbb{T}[r]}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
where $\eta_{\theta,v}^{l}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the set of wave packets of scale $r$. Since $(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$ is essentially supported in $\theta$ (Strictly speaking, this should be $C\theta$ for some constant $C$. However, it is convenient to omit the constant,as it does not impact the main arguments. Therefore, we will continue to do so in the following discussion.), we can choose bump functions $\psi_{\theta}^{\prime}$ so that $\psi_{\theta}^{\prime}=1$ on a $r^{-1/2}$ neighborhood of supp $\psi_{\theta}$, and decays rapidly outside this region. Finally, we define
$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
and then
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}+\mbox{RapDec}(R)\|f\|_{2}$ (2.1)
@@
$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}$
We say $f$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mbox{RapRec}(R)\|f\|_{2}$
[…]
$L^{2}$-OrthogonalityBy Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
@@
Let $r\geq 1$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
@@
Like in the case of paraboloid [11], there are also some local versions of the $L^{2}$-orthogonality for the cone, which are direct corollaries of Lemma 2.1:
@@
Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$,
@@
Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
@@
[…]
Spatial concentration For a wave packet $\left(\theta,v,l\right)$, the restriction of $Ef^{l}_{\theta,v}$ to the ball $B\left(0,r\right)$ is essentially supported on a tube. More precisely, let
$T^{l}_{\theta,v}:=N_{r^{\delta}}\left(P^{l}_{\theta,v}+rL\left(\theta\right)\right),\ \ rL\left(\theta\right):=\left\{tL\left(\theta\right):-r\leq t\leq r\right\}$
be a $r$-tube with long direction $L\left(\theta\right)$ and mini direction $M\left(\theta\right)$, here
$L\left(\theta\right):=\left(-{\xi}_{\theta},1\right),\ \ M\left(\theta\right):=\left({\xi}_{\theta},1\right)$
Figure 2: The spatial structure of a tube.
It is well-known that $Ef^{l}_{\theta,v}$ is essentially supported on $T^{l}_{\theta,v}$. Also, there is a standard $L^{2}$-estimate for $Ef$, see [6, 15] for details.
@@
For a function $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, we have
@@
Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
@@
[…]
Comparing wave packet decompositions at different scales Let $r^{1/2+2\delta}\leq\rho\leq r$ and $y\in B\left(0,r\right)$. Suppose $g$ is concentrated on the $r$-scale wave packet $\left(\theta,v,l\right)$ such that $T_{\theta,v,l}\cap B\left(y,\rho\right)\neq\emptyset$. To apply induction on scale, we need decompose a function into wave packet at different scales. We break up $g$ into wave packets over the ball $B\left(y,\rho\right)$ at scale $\rho$ as follows. First, we apply a transform to recenter $B\left(y,\rho\right)$ at the origin, i.e., $\tilde{g}\left(\xi\right):=e^{i{\psi}_{y}\left(\xi\right)}g\left(\xi\right)$, here
${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}\left|\xi\right|$
@@
${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{D}\mathrm{ist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{D}\mathrm{ist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\lesssim r^{\delta}\right)\right\}$
here
@@
and $P^{0}_{\theta}$ is the tine plates centered at origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction ${\xi}_{\theta}$.
@@
For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, there holds
@@
here $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
In [15] the relationship between the two wave packet decompositions was studied in detail. For example, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$, then which scale $\rho$ wave packets will $\tilde{g}$ be concentrated on? We state the results from [15] below.

### v7

# On Cone Restriction Estimates in Higher Dimensions

@@
Abstract.
We revisit the Ou-Wang’s approach to the cone restriction problem via polynomial partitioning. By recasting their inductive scheme as a recursive algorithm and incorporating the nested polynomial Wolff axioms, we obtain improved bounds for cone restriction estimates in higher dimensions.
1. Introduction
1.1. Statement of results
In this paper, we obtain a modest improvement for the cone restriction problem in higher dimensions. The Fourier restriction problem is one of the central open questions in harmonic analysis. It concerns the possibility of restricting the Fourier transform of an $L^{p}$ function to a curved hypersurface. Stein [16] conjectured that, for a well-curved hypersurface such as the sphere, the paraboloid, or the cone considered in this paper, the Fourier transform admits a bounded restriction operator from $L^{p^{\prime}}$ to $L^{q^{\prime}}$ for an appropriate range of exponents $p^{\prime}$ and $q^{\prime}$.
We now state the cone restriction conjecture precisely.
@@ 1.1. Statement of results
we have
$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{L^{p^{\prime}}(\mathbb{R}^{n})},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
@@
$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}+x_{n}|\xi|)}f(\xi)\,d\xi\$
where $B^{n-1}\subset\mathbb{R}^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that Conjecture 1.1 is equivalent to
@@
Let $n\geq 3$. For all $p,q$ satisfying (1.2), the estimates
$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p,q}\|f\|_{L^{q}(\mathbb{R}^{n-1})}$ (1.3)
holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
Conjecture 1.2 has been confirmed only in the cases $n=3,4,5$ , through the work of Barceló [1], Wolff [18], and Ou–Wang [15]. In higher dimensions, the best previously known range was obtained by Ou–Wang [15], whose results can be stated as follows.
@@
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n$ and
@@
then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
This theorem yields the following corollary.
Corollary 1.4 ([15], Theorem 1).
Let $n\geq 3$. For all $p$ satisfying
$p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$. (1.4)
the estimate
$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$ (1.5)
holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
We now state our main result. Given integers $n\geq 3$ and $k$ with $2\leq k\leq n-2$, define $p(n,k)$ and $q(n,k)$ by
$p(n,k):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
$\frac{1}{q(n,k)}=\frac{1}{p(n,k)}+\frac{1}{4n-8}\big(\frac{6}{p(n,k)}-1\big)$
For $q\geq 2$, we define $p_{n,k}(q)$ by
[…]
where $\alpha_{n,k}(q)$ is defined by
$\frac{1}{q}=\frac{\alpha_{n,k}(q)}{2}+\frac{1-\alpha_{n,k}(q)}{q(n,k)}$
@@
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n-2$ and
[…]
then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
Arguing as in [12], we obtain
@@
For $n\geq 3$ and all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$, the estimate (1.5) holds whenever
$p>2+\lambda n^{-1}+O(n^{-2}),\quad\lambda=2.596$… (1.6)
Observe that (1.4) implies that the estimate (1.5) holds for all $f:\mathbb{R}^{n-1}\to\mathbb{C}$ whenever
$p>2+\frac{8}{3n}+O(n^{-2}).$
Consequently, Corollary 1.6 yields an improved $L^{p}\to L^{p}$ bound for the cone restriction conjecture in higher dimensions.
Let $B_{R}\subset\mathbb{R}^{n}$ denote a ball of radius $R$. By a real interpolation of restricted-type, together with a standard $\epsilon$-removal argument (see [17, 4, 21]) and a bilinear interpolation (see [15]), Theorem 1.3 reduces to
Proposition 1.7.
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ satisfies $2\leq q\leq p$, $2\leq k\leq n-2$, and
$\begin{cases}p\geq p_{n,2}(q),q^{\prime}\leq\frac{n-2}{n}p&\text{if $k=2$},\\[4.0pt] p\geq p_{n,k}(q),p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\text{if $k\geq 3$}.\\[4.0pt] \end{cases},$
then the estimate
$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (1.7)
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
The Fourier restriction problem has been the focus of extensive study for decades. For hypersurfaces with non-vanishing Gaussian curvature, such as the sphere and paraboloid, restriction estimates are closely tied to several central problems in harmonic analysis, discrete geometry, and PDE, including the Bochner–Riesz conjecture, the Kakeya conjecture, and the Schrödinger maximal estimate. Furthermore, Fourier restriction has applications beyond harmonic analysis, impacting fields such as geometric measure theory (e.g., Falconer’s distance set conjecture) and number theory (e.g., counting solutions to Diophantine equations). However, sharp restriction estimates are known only for very few surfaces and dimensions. For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$. It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [14]), which highlights the potential applications of our results.
Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established an improved results for paraboloids. Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7]. In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, 25], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [22, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
In this article, we refine the results presented in [15] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [11, 12]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [15] as a recursive process. Compared to the paraboloid case addressed in [11, 12], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor.
1.2. Structure of the article
@@
1.3. Notational Conventions
@@
In this paper, we work with Schwartz functions $f$ and $g$ supported in $2\overline{B}^{\,n-1}\setminus B^{n-1}$. We write $x\in\mathbb{R}^{n}$ as $x=(x^{\prime},x_{n})$. The parameters $n$, $k$, and $\epsilon$ are fixed, and all implicit constants are allowed to depend on them.
We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ by
$\delta_{0}:=\epsilon^{10},\qquad\delta_{j}:=\delta_{j-1}^{10}\quad(1\leq j\leq n),\qquad\delta:=\delta_{n}^{10}.$
We write $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We write $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We write $A\ll B$ or $B\gg A$ if $CA\leq B$ for some admissible constant $C\geq 1$ that may be taken arbitrarily large. We write $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$.
Let #A\#A denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}\subset\mathbb{R}^{n}$ denote an arbitrary ball of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$.
We say that $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$. We write $A=\mathrm{RapDec}(r)$ if $A\lesssim_{N}r^{-N}$. We say that a function $F$ is essentially supported in $\Omega$ if
$F(x)\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}\quad\text{for all }x\notin\Omega.$
1.4. Acknowledgements
The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions. The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
1.5. Funding
This research received no external funding.
2. Preliminaries
2.1. Wave Packet Decomposition
@@
Wave packet decomposition is a central tool in our argument. We briefly recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
Definition. Fix $R\gg 1$ and let $R^{\delta_{0}}\leq r\leq R$. Cover $2\overline{B}^{\,n-1}\setminus B^{n-1}$ by finitely overlapping $r^{-1/2}$-sectors $\theta\in\Theta[r]$. A sector $\theta\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ is called an $r^{-1/2}$-sector if it has unit length in the radial direction and angular width $r^{-1/2}$. We denote by $\Theta[r]$ the collection of all such finitely overlapping $r^{-1/2}$-sectors. Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, we write
$f=\sum_{\theta\in\Theta[r]}f_{\theta},\qquad f_{\theta}:=f\,\psi_{\theta},$
where $\{\psi_{\theta}\}$ is a smooth partition of unity subordinate to the cover $\{\theta\}$. For each $\theta$, we further decompose $f_{\theta}$ as follows. First, decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length $r^{1/2+\delta}$ centered at $v\in r^{1/2+\delta}\mathbb{Z}^{n-1}$. Then
$\widehat{f_{\theta}}=\sum_{v}\widehat{f_{\theta}}\,\eta_{v},$
where $\{\eta_{v}\}$ is a smooth partition of unity subordinate to this cover.
Next, for each $v$, we decompose $\widehat{f_{\theta}}\eta_{v}$ into finer pieces adapted to $\theta$. Cover the cube centered at $v$ by parallel thin plates $P_{\theta,v}^{l}$ of radius $r^{1/2+\delta}$, thickness $r^{\delta/2}$, and normal direction $\xi_{\theta}$, where $\xi_{\theta}$ denotes the unit vector pointing in the central direction of the sector $\theta$. Hence,
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee},$
where $\{\eta_{\theta,v}^{l}\}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the collection of wave packets at scale $r$.
Since $\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}$ is essentially supported in $\theta$ (more precisely in $C\theta$ for some constant $C$), we choose bump functions $\psi_{\theta}^{\prime}$ such that $\psi_{\theta}^{\prime}=1$ on an $r^{-1/2}$–neighbourhood of $\operatorname{supp}\psi_{\theta}$ and decays rapidly outside this region. We then define
$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}.$
Consequently,
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}\;+\;\mathrm{RapDec}(R)\,\|f\|_{2}.$ (2.1)
@@
$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}.$
We say that $f$ is concentrated on the wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}.$
$L^{2}$-Orthogonality. By Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
@@
Suppose $r\geq 1$. For any $\mathcal{T}\subseteq\mathbb{T}[r]$, there holds
@@
As in the paraboloid case [11], there are corresponding local versions of the $L^{2}$-orthogonality for the cone, which follow directly from Lemma 2.1:
@@
Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$, we have
@@
Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$, we have
@@
Spatial Concentration. For a wave packet $(\theta,v,l)$, the restriction of $Ef_{\theta,v}^{l}$ to the ball $B(0,r)$ is essentially supported in a tube. More precisely, define
$T_{\theta,v}^{l}:=N_{r^{\delta}}\!\left(P_{\theta,v}^{l}+rL(\theta)\right),\qquad rL(\theta):=\{\,t\,L(\theta):-r\leq t\leq r\,\},$
which is an $r$-tube with long direction $L(\theta)$ and short direction $M(\theta)$, where
$L(\theta):=(-\xi_{\theta},\,1),\qquad M(\theta):=(\xi_{\theta},\,1).$
It is well known that $Ef_{\theta,v}^{l}$ is essentially supported in $T_{\theta,v}^{l}$.
2.2. A Standard $L^{2}$ Estimate.
Moreover, there is a standard $L^{2}$ estimate for EfEf; see [6, 15] for details.
@@
Suppose $f$ is a Schwartz function and supported in $2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, then
@@
Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on $R$-scale wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
@@
2.3. Comparing Wave Packet Decompositions at Different Scales.
Let $r^{1/2+2\delta}\leq\rho\leq r$ and let $y\in B(0,r)$. Suppose $g$ is concentrated on an $r$-scale wave packet $(\theta,v,l)$ with $T_{\theta,v,l}\cap B(y,\rho)\neq\emptyset$. To perform induction on scales, we decompose $g$ into wave packets at the smaller scale $\rho$ inside the ball $B(y,\rho)$. We begin by recentering $B(y,\rho)$ at the origin via the modulation
$\tilde{g}(\xi):=e^{i\psi_{y}(\xi)}\,g(\xi),$
where
${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}|\xi|$
@@
${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{Dist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{Dist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\right)\lesssim r^{\delta}\right\}$
where
@@
and $P^{0}_{\theta}$ denotes the thin plate centered at the origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction $\xi_{\theta}$.
@@
For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, we have
@@
where $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
The relationship between these two wave packet decompositions was studied in detail in [15]. In particular, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$, one may ask which wave packets at scale $\rho$ the function $\tilde{g}$ is concentrated on. We recall the relevant results from [15] below.